To find the height of a tower, a surveyor positions a transit that is 2 m tall at a spot 35 m from the base of the tower. She measures the angle of elevation to the top of the tower to be 58 degrees. What is the height of the tower, to the nearest meter?
Answer by pushpaharan(47)

if you draw a right triangle height will be opposite side, and the distance from the tower is adjacent.
use trigonometry
tan(A) = opposite/ adjacent
tan(58) = h/35
h = 35 * tan(58)
= 56 m
but actual height of the tower will be 56+2 = 58 m